# Multi-objective Optimization in CPU Design Space Exploration: Attention is All You Need

### v1

@@
Abstract.
Design space exploration (DSE) enables architects to systematically evaluate various design options, guiding decisions on the most suitable configurations to meet specific objectives such as optimizing performance, power, and area. However, the growing complexity of modern CPUs has dramatically increased the number of micro-architectural parameters and expanded the overall design space, making DSE more challenging and time-consuming. Existing DSE frameworks struggle in large-scale design spaces due to inaccurate models and limited insights into parameter impact, hindering efficient identification of optimal micro-architectures within tight timeframes.
In this work, we introduce AttentionDSE [^1] The code is now available on anonymous GitHub.. Its key idea is to use the attention mechanism to establish a direct mapping of micro-architectural parameters to their contributions to predicted performance. This approach enhances both the prediction accuracy and interpretability of the performance model. Furthermore, the weights are dynamically adjusted, enabling the model to respond to design changes and effectively pinpoint the key micro-architectural parameters/components responsible for performance bottlenecks. Thus, AttentionDSE accurately, purposefully, and rapidly discovers optimal designs. Experiments on SPEC 2017 demonstrate that AttentionDSE significantly reduces exploration time by over 80% and achieves 3.9% improvement in Pareto Hypervolume compared to state-of-the-art DSE frameworks while maintaining superior prediction accuracy and efficiency with an increasing number of parameters.
1. Introduction
Design space exploration (DSE) is an essential phase in CPU design, characterized by an iterative exploration that focuses on balancing performance, power, and area (PPA) (PPA1, ; PPA2, ; PPA3, ; PPA4, ). Typically, the DSE can be formulated as a Bayesian optimization (BO) process to find multi-objective trade-offs. A prevailing DSE framework primarily comprises two components: the surrogate model and the acquisition function. Specifically, the surrogate model predicts the PPA of micro-architectural parameters, bypassing time-consuming simulations (PPA2, ; blackbox2, ; blackbox3, ; BOOMExplorer, ; MoDSE, ), and the acquisition function identifies candidate optimal alternatives for multi-objective optimization, guiding the DSE iterations (nonblackbox1, ; nonblackbox2, ; nonblackbox3, ; nonblackbox4, ; ActBoost, ; bottleneck_analysis1, ).
Concurrently, the growing complexity of modern CPUs (PowerPC_cpu, ; AMD_cpu, ; superscalar_cpu, ) results in an exponential increase in micro-architectural parameters, known as the high-dimensional design space. High-dimensional DSE imposes significant limitations on current DSE frameworks, making it difficult to identify optimal designs within a reasonable timeframe, with some even failing to find optimal solutions altogether. Specifically, it presents three key challenges.
@@ 1. Introduction
    First, the effectiveness of DSE frameworks is constrained by the limited accuracy and speed of surrogate models in high-dimensional design spaces. To expedite DSE, various statistical regression methods (ActBoost, ; MoDSE, ; BO1, ; BO2, ; BOOMExplorer, ; Archranker, ) have been proposed as surrogate models to predict the performance metrics of micro-architecture. However, these models suffer from an exponential increase in training and inference time (random_forest, ; GPR, ) and their accuracy decreases with the expanding of the design space (dimgpr, ; dimencurse, ).
@@
    Second, the effectiveness of DSE frameworks is limited by the inefficiency of acquisition functions in high-dimensional design spaces. Acquisition functions are crucial in the DSE framework, as they identify the candidate optimal design for exploration. However, most acquisition functions depend heavily on expert knowledge to identify the most potential sample in the design space (ActBoost, ; MoDSE, ; BOOMExplorer, ), or on complex feature engineering (ArchExplorer, ) which requires a time-consuming process to identify the bottleneck of the current micro-architecture. Moreover, some of them even depend on exhaustive search (MoDSE, ; BOOMExplorer, ; ActBoost, ), which makes the acquisition functions impractical in high-dimensional design spaces.
@@
    Third, recent DSE frameworks overlook the impact of individual micro-architectural parameters on final performance metrics, missing valuable insights into their contributions. This issue primarily arises from the surrogate model’s inability to accurately capture the contribution of each parameter to overall performance (MoDSE, ; BOOMExplorer, ; ActBoost, ; PPA1, ; challenge3_3, ). Additionally, the separation of the surrogate model and acquisition function as independent entities in previous efforts further exacerbates this issue, as it impedes the seamless information exchange between them (ArchExplorer, ; MoDSE, ; BOOMExplorer, ; ActBoost, ).
In this work, we introduce AttentionDSE, a novel DSE framework that leverages the attention mechanism to accurately, purposefully, and rapidly identify optimal CPU designs in the high-dimensional design space. The key idea of AttentionDSE lies in its use of attention-based models to establish a direct mapping between micro-architectural parameters and performance metrics. This approach improves both the accuracy and interpretability of the performance model. Additionally, the attention weights are dynamically adjusted throughout the exploration process, allowing the model to adapt efficiently to design changes. This dynamic adjustment enables more effective identification of key micro-architectural parameters responsible for performance bottlenecks. As a result, AttentionDSE not only accelerates the design space exploration process but also ensures more precise and purposeful optimization, enabling rapid discovery of optimal micro-architectural designs.
Overall, our contributions are as follows:
@@
    We transform the DSE into a sequence prediction problem, introducing AttentionDSE that synergizes the surrogate model and acquisition function and leverages the attention mechanism for high-dimensional DSE on CPU.
@@
    We introduce a Perception-Driven Attention (PDA) mechanism, which utilizes micro-architectural insights to focus on relevant parameters, minimizing computational overhead of irrelevant parameters to reduce the training and inference time of attention-based performance model.
@@
    We propose an Attention-aware Bottleneck Analysis (ABA) algorithm, which uses attention weights to purposefully and quickly pinpoint bottlenecks, guiding the exploration.
@@
    Experimental results show that AttentionDSE reduces exploration time by over 80% and improves Pareto Hypervolume (PHV) by 3.9% compared to state-of-the-art (SOTA) DSE frameworks. Additionally, as the number of micro-architectural parameters increases, AttentionDSE keeps stable prediction accuracy and avoids significant increases in training and inference time.
2. Background
@@ 2.1. Attention Mechanism
(1) $\operatorname{Attention}(Q,K,V)=\operatorname{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V.$
It operates using three key vectors: query (Q), key (K), and value (V). The attention mechanism calculates a score by comparing the query and key vectors, which are normalized via a softmax function to produce attention weights. These weights then compute a weighted sum of the value vectors, emphasizing the most critical information. This process is especially effective in self-attention, where each input element attends to all others, and in multi-head attention, which captures different aspects of the input through multiple attention heads. Attention layers are foundational in models like transformers (transformer, ), enabling flexible and efficient learning of dependencies within sequences.
2.2. Design Space Exploration in CPU Design
DSE is a critical technique in the optimization and design process, primarily used to identify and analyze various design options within CPU design (PPA1, ; PPA2, ; PPA3, ; PPA4, ). Its goal is to systematically explore all possible design alternatives to find the optimal or near-optimal solutions that meet specific design requirements and constraints (ArchExplorer, ; MoDSE, ; BOOMExplorer, ; ActBoost, ; wangduo1, ; wangduo2, ).
To expedite the process, prevailing DSE frameworks adhere to the fundamental principles of BO (MoDSE, ; BOOMExplorer, ; ActBoost, ; Archranker, ). BO leverages a surrogate model that serves as a predictive tool, approximating the underlying objective function (GPR, ; random_forest, ; ensemble_learning, ). Additionally, an acquisition function drives the optimization process in DSE frameworks, efficiently navigating the DSE by identifying promising design points (PPA2, ), which is a set of micro-architectural parameters, for evaluation (ArchExplorer, ). The above two facilitate effective exploration of the design space within a reasonable time (BO1, ; BO2, ).
Figure 1. The workflow of prevailing DSE framework. ❶~❹ form the exploration loop.
Fig. 1 illustrates the core workflow of the BO-based DSE framework. Prior to the exploration stage, the framework first trains the surrogate model using sampling methods, such as Plackett-Burman ranking(PBrank, ; MoDSE, ), to generate the training set (①, ②, ③). Once the surrogate models are trained, the exploration stage begins. The acquisition function selects a design point from the design space (❶) and sends it to the surrogate models to evaluate the PPA (❷). The framework then iteratively updates the Pareto-optimal set and repeats this process (❸, ❹).
2.3. Pareto Optimal Set
@@
(2) $\forall i\in[1,n],f_{i}\left(x^{*}\right)\leq f_{i}(x);\exists j\in[1,n],f_{j}\left(x^{*}\right)<f_{j}(x).$
@@
(3) $\Omega=\left\{x\mid x^{*}\nprec x,\forall x^{*}\in D\right\}.$
@@
Figure 2. The Expansion of PHV: (a) original PHV; (b) design points excelling in performance or power extend the PHV.
3. Related Works
@@
For the early work, Mariani et al. (relative_work1, ) use a Kriging model (Kriging, ) with exponential correlation as a surrogate to predict the Pareto nondominated rank distribution. They then iteratively optimize using the Expected Improvement (EI) of non-dominated levels as the acquisition function. CASH (relative_work2, ) compares multiple machine learning models and finds that BO achieves superior exploration results, by using a random forest as the surrogate model and EI as the acquisition function. Wang et al. (relative_work3, ) employ an adaptive component selection and smoothing operator as the surrogate model and the expected Hypervolume Improvement (HVI) of the Pareto frontier as the acquisition function.
@@
For the blackbox DSE framework, ActBoost (ActBoost, ) integrates statistical methods and machine learning to prune the design space and determine the training set. For the surrogate model, ActBoost uses two AdaBoost.RT (AdaBoost.RT, ) models with active learning. For the acquisition function, it calculates the prediction uncertainty of each unsampled design point and selects one with the highest uncertainty. By combining statistical sampling and the boosting technique, ActBoost enables efficient and accurate DSE tasks. BOOMExplorer (BOOMExplorer, ) employs Gaussian Process Regression (GPR) (GPR, ) as the surrogate model and leverages EI (EI, ) as the acquisition function, searching for the next design point by analyzing data distribution characteristics. By embedding prior knowledge, BOOMExplorer achieves superior compared to analytical methods. MoDSE (MoDSE, ) leverages ensemble learning for more accurate prediction. It combines AdaBoost (AdaBoost, ) and Gradient Boosted Regression Trees (GBRT) (GBRT, ) as the surrogate model and uses HVI with a full-space search for the acquisition function. MoDSE introduces a uniformity-aware selection algorithm for efficient multi-objective DSE and a Pareto-rank-based sample weight generation algorithm to train the surrogate model, enhancing the DSE outcomes.
@@
For the DSE frameworks based on bottleneck analysis, ArchExplorer (ArchExplorer, ) adopts the dynamic event-dependence graph (DEG) (DEG, ) for critical path analysis and identifies bottlenecks in the current micro-architecture to guide the design of micro-architecture. It couples this approach with a software simulator (Gem5) (gem5, ) as the surrogate model. ArchExplorer focuses on refining the acquisition function to reduce the need for domain knowledge in mechanistic models. Explainable-DSE (Explainable_DSE, ) is the SOTA framework for DSE tasks in deep neural network (DNN) accelerators. By leveraging the structure and characteristics of neural networks, Explainable-DSE constructs a bottleneck model to identify design bottlenecks. It enables targeted optimizations to mitigate issues such as high latency and power consumption, while also providing explanations for the design adjustments made and their impact on performance. This framework is primarily used for the joint DSE of DNN accelerators and DNN algorithms.
4. Motivation
@@
4.1. Challenges in Current DSE Studies
Modern CPU design is becoming increasingly complex, with a growing number of micro-architectural parameters. The exploration in such a significantly large design space is known as the high-dimensional DSE. Current DSE frameworks face three primary challenges in high-dimensional tasks.
First, DSE frameworks struggle with effectiveness due to the limited accuracy and efficiency of surrogate models in high-dimensional design spaces. The surrogate models in prevailing DSE frameworks (ActBoost, ; MoDSE, ; BOOMExplorer, ; Archranker, ), which typically use statistical regression methods, are simple while vulnerable. Specifically, prediction accuracy can decline as dimensionality increases. For instance, when the number of parameters exceeds 75, BOOMExplorer struggles to accurately predict and associate micro-architectural parameters with performance metrics.
This issue arises from the sparser feature space in high-dimensional condition (dimencurse, ), indicating the need for more training data to accurately capture the data distribution. However, the training and inference time increases greatly as the size of the training set grows. For instance, with a training set of size $n$, the complexity of training the inverse of the covariance matrix typically reaches $O(n^{3})$ and the inference through the inverse of the covariance matrix is $O(n^{2})$. This considerable computational cost renders high-dimensional design spaces impractical.
Second, the performance of DSE frameworks is hindered by the inefficiency of acquisition functions when dealing with high-dimensional design spaces. Acquisition functions are crucial for finding optimal design configurations, but they face challenges in high-dimensional design spaces. Current methods often rely on expert knowledge (ActBoost, ; MoDSE, ; BOOMExplorer, ) or complex feature engineering (ArchExplorer, ) to identify the bottleneck of the current micro-architecture, which can introduce biases, require significant time, and lack scalability. For instance, ArchExplorer only analyses the first hundred thousand of the instructions and achieves the optimal architecture with the number of parameters 21 over 15 days (ArchExplorer, ). Moreover, some methods rely on large-scale or even exhaustive search (MoDSE, ; ActBoost, ; BOOMExplorer, ). In high-dimensional spaces, this exhaustive search becomes computationally prohibitive due to the enlarged design space. For example, MoDSE can only perform with the number of parameters 10 (totaling 36,864 design points) (MoDSE, ). When the number of parameters increases to 30, it may take $1.9\times 10^{29}$ years. The high computational cost of such searches slows down the DSE process, delaying the identification of high-performance designs. This inefficiency slows down the DSE process, highlighting the need for more advanced and efficient acquisition strategies to handle complex, high-dimensional problems.
Third, conventional DSE frameworks fail to recognize the impact of each parameter on performance metrics. This limitation can cause these frameworks to overlook critical insights into how individual parameters impact the performance metrics, making it difficult to quickly and accurately identify bottlenecks in the current design. This phenomenon can be attributed to two primary reasons. Firstly, existing surrogate models cannot unveil the contributions of each parameter to the performance metrics because they rely on mathematical analysis (ActBoost, ; Archranker, ) or simplistic black-box models (BOOMExplorer, ; MoDSE, ) for prediction. Secondly, prevailing DSE frameworks typically adhere to a two-stage DSE scheme, treating the surrogate model and acquisition function as independent entities. This separation inhibits the exchange of information, preventing a comprehensive understanding of parameter efficacy (BOOMExplorer, ; MoDSE, ; ActBoost, ; ArchExplorer, ). Moreover, the lack of integration between these two components hinders the ability of the acquisition function to adaptively refine the exploration results based on feedback from the surrogate model, resulting in suboptimal performance and less effective DSE.
4.2. Opportunity of Attention in DSE Tasks
DSE is a crucial process in CPU design, often consisting of two important tasks: prediction and bottleneck analysis. Prediction helps estimate the performance of various design configurations, while bottleneck analysis identifies critical constraints that limit system efficiency. Both tasks are essential to guide CPU design effectively.
Recent advances in attention mechanisms offer significant advantages in addressing these challenges, proving to be highly effective for both performance prediction and bottleneck analysis. First, attention has been demonstrated to excel in regression tasks (attention_regression1, ; attention_regression2, ), accurately capturing complex relationships between inputs and outputs. Second, attention mechanisms inherently highlight the importance of different parameters by assigning varying weights, thus enabling a more fine-grained analysis of key design factors. These strengths align directly with the core needs of DSE, where both accurate performance prediction and the identification of critical design parameters are essential for driving effective design exploration.
5. Design of AttentionDSE

### v2

@@
Abstract
Design Space Exploration (DSE) is essential to modern CPU design, yet current frameworks struggle to scale and generalize in high-dimensional architectural spaces. As the dimensionality of design spaces continues to grow, existing DSE frameworks face three fundamental challenges: (1) reduced accuracy and poor scalability of surrogate models in large design spaces; (2) inefficient acquisition guided by hand-crafted heuristics or exhaustive search; (3) limited interpretability, making it hard to pinpoint architectural bottlenecks.
In this work, we present AttentionDSE, the first end-to-end DSE framework that natively integrates performance prediction and design guidance through an attention-based neural architecture. Unlike traditional DSE workflows that separate surrogate modeling from acquisition and rely heavily on hand-crafted heuristics, AttentionDSE establishes a unified, learning-driven optimization loop, in which attention weights serve a dual role: enabling accurate performance estimation and simultaneously exposing the performance bottleneck. This paradigm shift elevates attention from a passive representation mechanism to an active, interpretable driver of design decision-making.
Key innovations include: (1) a Perception-Driven Attention mechanism that exploits architectural hierarchy and locality, scaling attention complexity from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n)$ via sliding windows; (2) an Attention-aware Bottleneck Analysis that automatically surfaces critical parameters for targeted optimization, eliminating the need for domain-specific heuristics.
Evaluated on high-dimensional CPU design space using the SPEC CPU2017 benchmark suite, AttentionDSE achieves up to 3.9% higher Pareto Hypervolume and over 80% reduction in exploration time compared to state-of-the-art baselines.
footnotetext: Corresponding author is Mingyu Yan (yanmingyu@ict.ac.cn).
I Introduction
As modern CPUs continue to integrate more aggressive micro-architectural features and support increasingly diverse workloads, design space exploration (DSE) has become a crucial phase in the processor design cycle. DSE seeks to identify optimal configurations across a vast space of micro-architectural parameters under multiple conflicting objectives, such as performance, power, and area (PPA) [21, 1, 12, 31]. To accelerate this process, prior work often formulates DSE as a multi-objective Bayesian Optimization (BO) task, leveraging a surrogate model to approximate PPA [1, 6, 9, 3, 42] and an acquisition function to guide design refinement [18, 19, 29, 38, 22, 28].
However, with the escalating complexity of CPU micro-architecture, the dimensionality of the design space is growing rapidly—reaching hundreds of tunable parameters [26, 27, 20]. This high-dimensional design space presents three critical challenges for existing DSE frameworks:
@@ I Introduction
    Inaccuracy and poor scalability of surrogate models: Existing statistical models (e.g., GPR, ensemble trees) struggle to maintain accuracy and scalability in high-dimensional spaces [22, 42, 16, 34, 8, 14], suffering from sparsity of data and a long training time [5, 45].
@@
    Ineffective acquisition functions: Most acquisition strategies rely on heuristic rules [42, 3], feature engineering [2], or exhaustive search [42, 3, 22]—which become impractical as the micro-architectural parameter count increases.
@@
    Lack of parameter-wise interpretability: Prior work offers limited visibility into the performance sensitivity or bottleneck contribution of each micro-architectural parameters [42, 3, 22], impeding informed design decisions.
Opportunity. We observe that the above challenges are fundamentally rooted in the limited expressiveness and decoupled nature of existing DSE frameworks, where the surrogate model and acquisition function are treated as separate black boxes.
Meanwhile, recent advances in attention-based deep learning—particularly transformers—offer a compelling alternative. First, attention mechanisms excel at modeling long-range dependencies across high-dimensional inputs, enabling them to accurately learn how diverse micro-architectural parameters jointly influence PPA. Second, the modular and parallelizable nature of attention layers ensures scalability as the number of micro-architectural parameters increases. Third, attention naturally learns to weight the influence of individual parameters when predicting target performance metrics, providing built-in interpretability that is essential for identifying architectural bottlenecks and guiding targeted design refinements.
These properties make attention-based models uniquely positioned to address the key challenges of DSE. By unifying surrogate prediction and acquisition within an attention-driven neural architecture, it becomes possible to achieve accurate, scalable, and explainable exploration—without relying on hand-crafted features or domain-specific rules. This convergence of modeling power, transparency, and scalability presents a timely opportunity to rethink DSE methodology for the era of high-dimensional CPU design.
Our Proposal. To address the limitations of existing frameworks in high-dimensional CPU DSE, we propose AttentionDSE—an attention-driven framework that enables end-to-end, interpretable, and scalable DSE. AttentionDSE unifies performance prediction and exploration guidance within a single transformer-based neural architecture. Its predictor captures both the performance sensitivity and bottleneck contribution of each micro-architectural parameter, with attention weights serving a dual role: effectively supporting accurate PPA estimation and naturally exposing parameter-wise bottlenecks to directly guide the optimization process in a data-driven and interpretable manner.
To further enhance scalability and interpretability in high-dimensional settings, two key innovations are introduced:
@@
    Perception-Driven Attention (PDA): A hardware-aware serialization and localized attention mechanism that captures architectural hierarchy and local connectivity, reducing attention complexity from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n)$ while greatly improving scalability without loss of modeling fidelity.
@@
    Attention-aware Bottleneck Analysis (ABA): An exploration algorithm that interprets the attention weights to automatically identify performance-critical parameters, enabling fine-grained and explainable design updates without relying on expert knowledge or external heuristics.
Together, these components enable AttentionDSE to form a closed-loop, data-driven DSE workflow that supports scalable, interpretable, and automated exploration in high-dimensional design spaces. This accelerates convergence and yields actionable insights into micro-architectural trade-offs—paving the way for more intelligent and insight-driven architecture design.
We evaluate AttentionDSE on the SPEC CPU 2017 benchmark suite using an OoO CPU with high-dimensional design space. Compared to state-of-the-art DSE frameworks (e.g., BOOMExplorer [3], MoDSE [42], ArchExplorer [2]), AttentionDSE reduces exploration time by over 80%, achieves up to 3.9% improvement in Pareto Hypervolume (PHV), and maintains stable accuracy even in high-dimensional settings.
Contributions. This work makes the following contributions:
@@
    We propose AttentionDSE, the first end-to-end attention-driven DSE framework that unifies performance prediction and exploration within a single interpretable architecture.
@@
    We introduce PDA, a hardware-aware serialization and sliding attention mechanism that improves scalability in high-dimensional design spaces.
  • $\bullet$
    We develop ABA, a novel method that uses attention weights for interpretable, automated performance bottleneck analysis and design refinement.
  • $\bullet$
    We demonstrate significant improvements in accuracy, efficiency, and scalability over state-of-the-art DSE baselines on realistic CPU benchmarks.
II Background
@@ II-A Attention Mechanism
$\operatorname{Attention}(Q,K,V)=\operatorname{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V.$ (1)
It operates using three key vectors: query (Q), key (K), and value (V). The attention mechanism calculates a score by comparing the query and key vectors, which are normalized via a softmax function to produce attention weights. These weights then compute a weighted sum of the value vectors, emphasizing the most critical information. This process is especially effective in self-attention, where each input element attends to all others, and in multi-head attention, which captures different aspects of the input through multiple attention heads. Attention layers are foundational in models like transformers [40], enabling flexible and efficient learning of dependencies within sequences.
II-B Design Space Exploration in CPU Design
DSE is a critical technique in the optimization and design process, primarily used to identify and analyze various design options within CPU design [21, 1, 12, 31]. Its goal is to systematically explore all possible design alternatives to find the optimal or near-optimal solutions that meet specific design requirements and constraints [2, 42, 3, 22, 43, 41].
To expedite the process, prevailing DSE frameworks adhere to the fundamental principles of BO [42, 3, 22, 10]. BO leverages a surrogate model that serves as a predictive tool, approximating the underlying objective function [45, 5, 13]. Additionally, an acquisition function drives the optimization process in DSE frameworks, efficiently navigating the DSE by identifying promising design points [1], which is a set of micro-architectural parameters, for evaluation [2]. The above two facilitate effective exploration of the design space within a reasonable time [16, 34].
Figure 1: The prevailing DSE workflow. ❶˜❹ form the exploration loop.
Fig. 1 illustrates the core workflow of the BO-based DSE framework. Prior to the exploration stage, the framework first trains the surrogate model using sampling methods, such as Plackett-Burman ranking[39, 42], to generate the training set (①, ②, ③). Once the surrogate models are trained, the exploration stage begins. The acquisition function selects a design point from the design space (❶) and sends it to the surrogate models to evaluate the PPA (❷). The framework then iteratively updates the Pareto-optimal set and repeats this process (❸, ❹).
II-C Pareto Optimal Set
@@
$\forall i\in[1,n],f_{i}\left(x^{*}\right)\leq f_{i}(x);\exists j\in[1,n],f_{j}\left(x^{*}\right)<f_{j}(x).$ (2)
@@
$\Omega=\left\{x\mid x^{*}\nprec x,\forall x^{*}\in D\right\}.$ (3)
@@
Figure 2: The improvement of PHV: (a) original PHV; (b) design points excelling in performance or power extend the PHV.
III Related Works
Many efforts [10, 35, 44, 24, 3, 42, 2, 25, 22] accomplish the DSE tasks.
@@ III Related Works
For the early work, Mariani et al. [24] use a Kriging model [32] with exponential correlation as a surrogate to predict the Pareto nondominated rank distribution. They then iteratively optimize using the Expected Improvement (EI) of non-dominated levels as the acquisition function. CASH [35] compares multiple machine learning models and finds that BO achieves superior exploration results, by using a random forest as the surrogate model and EI as the acquisition function. Wang et al. [44] employ an adaptive component selection and smoothing operator as the surrogate model and the expected Hypervolume Improvement (HVI) of the Pareto frontier as the acquisition function.
@@
For the blackbox DSE framework, ActBoost [22] integrates statistical methods and machine learning to prune the design space and determine the training set. For the surrogate model, ActBoost uses two AdaBoost.RT [36] models with active learning. For the acquisition function, it calculates the prediction uncertainty of each unsampled design point and selects one with the highest uncertainty. By combining statistical sampling and the boosting technique, ActBoost enables efficient and accurate DSE tasks. BOOMExplorer [3] employs Gaussian Process Regression (GPR) [45] as the surrogate model and leverages EI [46] as the acquisition function, searching for the next design point by analyzing data distribution characteristics. By embedding prior knowledge, BOOMExplorer achieves superior compared to analytical methods. MoDSE [42] leverages ensemble learning for more accurate prediction. It combines AdaBoost [33] and Gradient Boosted Regression Trees (GBRT) [30] as the surrogate model and uses HVI with a full-space search for the acquisition function. MoDSE introduces a uniformity-aware selection algorithm for efficient multi-objective DSE and a Pareto-rank-based sample weight generation algorithm to train the surrogate model, enhancing the DSE outcomes.
@@
For the DSE frameworks based on bottleneck analysis, ArchExplorer [2] adopts the dynamic event-dependence graph (DEG) [15] for critical path analysis and identifies bottlenecks in the current micro-architecture to guide the design of micro-architecture. It couples this approach with a software simulator (Gem5) [4] as the surrogate model. ArchExplorer focuses on refining the acquisition function to reduce the need for domain knowledge in mechanistic models. Explainable-DSE [11] is the SOTA framework for DSE tasks in deep neural network (DNN) accelerators. By leveraging the structure and characteristics of neural networks, Explainable-DSE constructs a bottleneck model to identify design bottlenecks. It enables targeted optimizations to mitigate issues such as high latency and power consumption, while also providing explanations for the design adjustments made and their impact on performance. This framework is primarily used for the joint DSE of DNN accelerators and DNN algorithms.
IV Motivation
@@
IV-A Challenges in Current DSE Studies
As modern CPUs continue to integrate increasingly aggressive micro-architectural features and support a wider range of workloads, their design complexity grows substantially. This complexity is reflected in the expanding number of tunable micro-architectural parameters—spanning pipeline width, core frequency, and so on. As a result, the design space has become significantly larger and more intricate, giving rise to the challenge of high-dimensional DSE.
Existing DSE frameworks, however, struggle to scale in such settings and face three fundamental challenges in high-dimensional exploration tasks.
First, the effectiveness of DSE frameworks is limited by the low accuracy and poor scalability of surrogate models in high-dimensional CPU design spaces. Most existing approaches [22, 42, 3, 10] use statistical models—such as linear regression or Gaussian processes—to predict PPA. While efficient on small search spaces, these models quickly break down as the number of design parameters grows. For instance, when the design space exceeds 75 parameters, BOOMExplorer struggles to maintain prediction accuracy.
This is mainly due to the “curse of dimensionality”: as the number of parameters increases, the design space becomes exponentially sparser, requiring significantly more samples to train reliable models [14]. Unfortunately, increasing training size leads to excessive overhead. For example, in Gaussian Process-based models, training requires matrix inversion with $O(n^{3})$ complexity and inference scales as $O(n^{2})$, where $n$ is the number of training samples. This makes it impractical to scale traditional surrogate modeling to realistic, high-dimensional CPU design scenarios.
Second, the performance of DSE frameworks is hindered by the inefficiency of acquisition functions in high-dimensional design spaces. Acquisition functions are critical for guiding exploration toward optimal micro-architectural configurations. However, in large parameter spaces, existing strategies become inefficient and poorly scalable. Many methods depend on domain-specific heuristics [22, 42, 3] or handcrafted features [2] to identify performance bottlenecks. These approaches are often biased, time-consuming, and difficult to generalize across architectures and workloads.
For example, ArchExplorer [2] analyzes only the first 100,000 instructions and takes 15 days to optimize an architecture with just 21 parameters. Some frameworks resort to large-scale or even exhaustive search [42, 22], which becomes computationally infeasible as dimensionality increases. MoDSE [42], for instance, operates over a 10-parameter space with 36,864 design points, but scaling to 30 parameters would require an astronomical $1.9\times 10^{29}$ years of search time.
These limitations highlight the need for acquisition strategies that can operate efficiently in high-dimensional settings—without relying on heuristics or exhaustive enumeration—while still identifying high-quality designs.
Third, existing DSE frameworks lack interpretability, making it difficult to reason about trade-offs and locate architectural bottlenecks. Conventional frameworks offer little insight into how individual micro-architectural parameters affect PPA. This makes it challenging to understand why a given design performs well or poorly, thereby hampering informed, targeted design refinements.
This problem stems from two core limitations. First, most surrogate models rely on either mathematical analysis [22, 10] or black-box predictors [3, 42], which fail to expose parameter-wise contributions to performance. Second, the standard two-stage DSE workflow treats performance prediction and design exploration as separate steps [3, 42, 22, 2], preventing feedback from the predictor from guiding future exploration. As a result, it is difficult for the acquisition function to adaptively prioritize parameters that are more performance-sensitive.
Without a unified, feedback-driven approach, existing methods are not only harder to interpret, but also less effective at navigating high-dimensional design spaces-especially when timely, bottleneck-aware decisions are required.
IV-B Why Attention for DSE?
The attention mechanism provides a powerful and well-suited solution to above challenges for several reasons:
  • $\bullet$
    Context-aware Interaction Modeling: Attention dynamically computes the relative importance of each parameter with respect to others in a given design configuration [7]. This enables the model to capture complex, conditional dependencies—such as how the benefit of a wide fetch stage may depend on the reorder buffer size or instruction window depth—without manual feature engineering.
  • $\bullet$
    Scalability and Structural Flexibility: With recent advances such as sparse attention, attention-based models can scale to hundreds of parameters while maintaining modeling fidelity. Moreover, attention can be adapted to reflect architectural hierarchies and locality, enabling better alignment with the structure of CPU design spaces.
  • $\bullet$
    Unified Modeling for Prediction and Guidance: Attention can bridge surrogate modeling and exploration by using a single predictor to both estimate PPA and inform the search direction. This eliminates the need for hand-crafted heuristics or separate analysis stages and enabling a more efficient, closed-loop DSE process.
  • $\bullet$
    Interpretability through Attention Weights: Unlike black-box predictors, attention mechanisms expose internal weights that reveal which parameters contribute most to performance metrics. This provides a natural interface for bottleneck analysis and informed refinement, supporting more targeted and efficient exploration.
These properties position attention not merely as a prediction tool, but as a unified interface between surrogate modeling and acquisition strategy. When properly integrated, attention mechanisms can enable an end-to-end, self-driven DSE framework that is both interpretable and scalable—paving the way for a new class of intelligent architecture design tools.
V Design of AttentionDSE